# Typical thermalization of low-entanglement states

### v3

@@ Abstract
A drawback of many of these approaches, in the case of both the ETH and typicality, is that they often require the assumption that the system is prepared in an initial state which is concentrated around a given energy $E$. While all states with sub-extensive energy variance are concentrated around their average energy, it is often not clear how strong the concentration needs to be in order for thermalization to be guaranteed given either a formulation of ETH or a typicality transformation. One often assumes that the support of the state vanishes outside of a narrow micro-canonical energy window; depending on how narrow this window is such a state generally cannot be efficiently prepared. Crucially, this excludes important classes of physical and operationally meaningful initial states, such as product states and states with low entanglement. These states, in addition to being fundamental to the study of lattice systems, are commonly the only accessible initial states to thermalization experiments in quantum simulators or numerical examination of thermalization, for example in the context of probing many-body localization [23, 24, 25, 26, 27], and are often considered in works on relaxation and equilibration [28, 29, 30].
@@
In this work, we provide this important missing piece of the puzzle by proving thermalization of low-entanglement states. We address this by studying states in an analog of the micro-canonical ensemble which is allowed to have support on the whole energy spectrum. Building upon and extending the work of Ref. [31], we show that, under precise and physically meaningful conditions, these states are locally equivalent to Gibbs states. We later show, generalizing the seminal work of Ref. [21], that such states crucially arise from the long-time dynamics of low-entanglement initial states under typical Hamiltonians. Our typicality transformation consists of a unitary operation that randomly mixes eigenstates with nearby energies. Related transformations have been introduced in Ref. [32] to compute higher order correlation functions under a generalized version of the ETH stemming from local (in energy) unitary invariance, and this has been connected to the theory of free probability in Ref. [33]. In our work, we explicitly demonstrate how such a transformation leads to equilibrium states being locally indistinguishable from thermal states. Moreover, by bounding the energy difference between eigenstates that get mixed, we elucidate how the original Hamiltonian is related to the transformed one. More specifically, given a local Hamiltonian $H$ and an initial low-entanglement state $\rho$, we define an ensemble of random Hamiltonians such that, if $H^{\prime}$ is a Hamiltonian drawn from this ensemble: 1. the Gibbs states of $H$ and $H^{\prime}$ are locally indistinguishable, and 2. with overwhelming probability the equilibrium state resulting from the unitary evolution $\mathrm{e}^{-\mathrm{i}H^{\prime}t}\rho\mathrm{e}^{\mathrm{i}H^{\prime}t}$ is locally indistinguishable from said Gibbs state. In addition, we give an assumption under which the dynamics of $H$ and $H^{\prime}$ are indistinguishable in the short-time regime. This new kind of Hamiltonian-typicality approach is rooted in physical grounds: it is unrealistic to have a system specified to infinite precision. This approach relaxes and overcomes the common idea that thermalization has to be proven for all Hamiltonians, but rather, it turns out to be typical for an overwhelming portion of meaningful settings.
@@
with $H^{\prime}=UHU^{\dagger}$. This means that the dynamics under $U$ and $UHU^{\dagger}$ are indistinguishable up to a time $t^{*}\sim N^{\kappa-\frac{1-\alpha}{D+1}}$. If $\kappa>0$, that is the BE error decays faster than the worst-case scenario, $\alpha$ can be chosen such that $t^{*}$ increases with the system size. In other words $H$ and $UHU^{\dagger}$ generate nearly the same dynamics for a time $t^{*}\sim\mathrm{poly}(N)$. Finally, we would like to emphasize that our rigorous approach allows to put on precise and solid ground some of the results obtained on equilibration in Ref. [43].
It is now worth noting that if both $\rho$ and $\sigma$ are translation-invariant, the averaging over regions in the definition of $D_{l}(\rho,\sigma)$ can be dropped, making the indistinguishability statement valid for any observable supported on an individual small region. We show that if the original Hamiltonian is translation-invariant, we recover this property to some extent in the equilibrium state of the perturbed Hamiltonian. More specifically, consider an observable $A$ supported in $C\in\mathcal{C}_{l}$ and $H$ translation-invariant. For $U$ drawn from $\mathcal{E}(\delta)$ we show that, with probability at least $1-\sum_{k}\exp(-c{d_{k}\epsilon}/{\|A\|^{2}_{\infty}})$ for a constant $c$, it holds, with $d_{k}=\mathrm{dim}(\mathcal{W}_{k})$,
$\left|\mathrm{tr}\left(\left(g_{\beta}(H)-\rho_{\infty}^{UHU^{\dagger}}\right)A\right)\right|\leq\epsilon+D_{l}\left(\rho_{\infty}^{UHU^{\dagger}},g_{\beta}(H)\right).$ (17)
Details and proofs are available in the Supplemental Material, Section IV. Notice that the probability is large assuming that the number of eigenstates in each windows grows with the system size. By applying this to $\rho=g_{\beta}(UHU^{\dagger})$ we also get translation invariance in the same sense for the randomized Gibbs state, that is, $\rho_{\infty}^{UHU^{\dagger}}$ can be replaced by $g_{\beta}(UHU^{\dagger})$ in Eq. 17.
@@
Turning to notions of dynamical thermalization, we have investigated the typical time-evolution of the expectation value of a generic observable $A$, $\langle A\rangle_{\rho}:=\mathrm{tr}\left(A\rho\right)$, under the evolution generated by $UHU^{\dagger}$. In the Supplemental Material, Section III, we show that, with probability at least $1-\sum_{k}\exp\left(-c{d_{k}\epsilon}/{\|A\|^{2}_{\infty}}\right)$ for a constant $c$, the time-evolution is bounded by
$|\langle A\rangle_{\rho^{UHU^{\dagger}}(t)}-\langle A\rangle_{\rho_{\infty}^{UHU^{\dagger}}}|\leq\epsilon+R(t),$ (18)
@@
Several natural open questions remain. Importantly, the problem of identifying precise properties of the Hamiltonian achieving a lower than worst case BE error. For these Hamiltonians, our result predicts that a vast ensemble of other Hamiltonians exists that all have indistinguishable short time dynamics and which eventually thermalize. Furthermore, while our randomization achieves thermalization, more work may be needed about exploring the precise physical mechanisms that can be held responsible for them. On the one hand, being able to implement such transformations in a controlled way might have implications in the form of algorithms for Gibbs state preparation. On the other hand, understanding under what condition an uncertainty in either state or Hamiltonian preparation can translate into a random energy-preserving perturbation such as the one we defined might shed light on the stability of many-body localization under realistic conditions. It is the hope that this work contributes to understanding that under a “trembling hand”, most systems follow the laws of quantum statistical mechanics.

### v4

@@ Abstract
A make or break question is left open, however: Why would one generically expect – starting with natural initial states such as low-entanglement states – quantum systems to relax to states that are operationally indistinguishable from Gibbs states by local measurements. This is the actual empirical observation of apparent thermalization that needs to be answered by a theoretical framework. The core prejudice to be overcome is not to prove thermalization for all Hamiltonians, but only for most such Hamiltonians that are slightly randomized, and still yield extremely similar dynamical predictions compared to the original Hamiltonian.
In this work, we provide an important missing piece of the puzzle of proving thermalization of low-entanglement states. We show that initial low-entanglement states thermalize under appropriately weakly randomized local Hamiltonians. In contrast to previous approaches [20, 21], we focus on making the randomization as weak as possible while still ensuring thermalization under the randomized Hamiltonian. Under a meaningful spectral smoothness assumption, the randomization can be chosen to be weak enough such that it does not affect the short-time dynamics. In addition, we show that the randomization does not affect the thermal properties of the Hamiltonian. With these last two points, we take steps towards bridging the gap between the spatially local description of the system given by a local Hamiltonian, and the energy-local description given by random matrix theory necessary to explain thermalization. Accepting that the Hamiltonian will anyway only be known to a finite precision, such a perturbative typicality argument goes a long way towards explaining why a dynamical convergence to Gibbs states is so ubiquitous in nature.
It is known that quantum lattice models that are prepared in typical states that are restricted to a narrow window centered around a given energy – invoking notions of equivalence of ensembles – will give rise to local expectation values that are indistinguishable from thermal ones [23]. While this is conceptually extremely insightful, it is not clear how such states can arise from physical dynamics of natural initial states. What is, in contrast, very natural is to consider low-entanglement initial conditions and natural time evolution. Such low-entanglement states, in addition to being fundamental to the study of lattice systems, are commonly the only accessible initial states to thermalization experiments in quantum simulators or numerical examination of thermalization, for example, in the context of probing many-body localization [24, 25, 26, 27, 28], and are often considered in works on relaxation and equilibration [29, 30, 31].
Concretely, given a local Hamiltonian $H$ and an initial low-entanglement state $\rho$, we define an ensemble of random Hamiltonians such that, if $H^{\prime}$ is a Hamiltonian drawn from this ensemble: 1. the Gibbs states of $H$ and $H^{\prime}$ are locally indistinguishable, and 2. with overwhelming probability the equilibrium state resulting from the unitary evolution $\mathrm{e}^{-\mathrm{i}H^{\prime}t}\rho\mathrm{e}^{\mathrm{i}H^{\prime}t}$ is locally indistinguishable from said Gibbs state. Importantly, we further give an assumption under which the dynamics of $H$ and $H^{\prime}$ are indistinguishable in the short-time regime. (See Fig. 1).
Figure 1: Illustration of the dynamics of the original and perturbed Hamiltonians. The distance of the new Hamiltonian from the original one is bounded by a parameter $\delta$. The plot shows how the evolution of an observable under $H$ or $H^{\prime}$ behaves: the two evolutions agree up to a time $\sim 1/\delta$, and the perturbed Hamiltonian equilibrates to a thermal value.
@@
We address this by studying states in an analog of the micro-canonical ensemble which is allowed to have support on the whole energy spectrum. Building upon and extending the work of Ref. [23], we show that, under precise and physically meaningful conditions, these states are locally equivalent to Gibbs states. We later show that such states arise from the long-time dynamics of low-entanglement initial states under typical Hamiltonians. Our typicality transformation, defining the ensemble of random Hamiltonians, consists of a unitary operation that randomly mixes eigenstates with nearby energies. Related transformations have been introduced in Ref. [32] to compute higher order correlation functions under a generalized version of the ETH stemming from local (in energy) unitary invariance, and this has been connected to the theory of free probability in Ref. [33]. In our work, we explicitly demonstrate how such a transformation leads to equilibrium states being locally indistinguishable from thermal states. Moreover, by elucidating how the original Hamiltonian is related to the transformed one, we motivate a meaningful, operational interpretation of our Hamiltonian-typicality approach: Given a physical system in the right conditions, one can find a local Hamiltonian $H$ which well describes the dynamics. Nevertheless, there is a host of other models $H^{\prime}$ which, if $H$ is well behaved, describe very similar dynamics as well as equilibrium properties, and are close to $H$. But $H$ cannot be known to infinite precision. This operational uncertainty can be modeled by our Hamiltonian ensemble.
@@
with $H^{\prime}=UHU^{\dagger}$. This means that the dynamics under $H$ and $UHU^{\dagger}$ are indistinguishable up to a time $t^{*}\sim N^{\kappa-\frac{1-\alpha}{D+1}}$. If $\kappa>0$, that is the BE error decays faster than the worst-case scenario, $\alpha$ can be chosen such that $t^{*}$ increases with the system size. In other words $H$ and $UHU^{\dagger}$ generate nearly the same dynamics for a time $t^{*}\sim\mathrm{poly}(N)$. Finally, we would like to emphasize that our rigorous approach allows to put on precise and solid ground some of the results obtained on equilibration in Ref. [43].
It is now worth noting that if both $\rho$ and $\sigma$ are translation-invariant, the averaging over regions in the definition of $D_{l}(\rho,\sigma)$ can be dropped, making the indistinguishability statement valid for any observable supported on an individual small region. We show that if the original Hamiltonian is translation-invariant, we recover this property to some extent in the equilibrium state of the perturbed Hamiltonian. More specifically, consider an observable $A$ supported in $C\in\mathcal{C}_{l}$ and $H$ translation-invariant. For $U$ drawn from $\mathcal{E}(\delta)$ we show, assuming all windows centered around an extensive energy contain exponentially many eigenvalues, that except with probability $\mathrm{e}^{-\Omega(N)}$ ,
$\bigg{|}\mathrm{tr}\bigg{(}\big{(}g_{\beta}(H)-\rho_{\infty}$ )${}^{UHU^{\dagger}}\big{)}A\bigg{)}\bigg{|}\leq$ (17)
$\mathrm{e}^{-\Omega(N)}+D_{l}\left(\rho_{\infty}^{UHU^{\dagger}},g_{\beta}(H)\right).$
Details and proofs are available in the Supplemental Material, Section V. By applying this to $\rho=g_{\beta}(UHU^{\dagger})$ we also get translation invariance in the same sense for the randomized Gibbs state, that is, $\rho_{\infty}^{UHU^{\dagger}}$ can be replaced by $g_{\beta}(UHU^{\dagger})$ in Eq. 17.
@@
Turning to notions of dynamical thermalization, we have investigated the typical time-evolution of the expectation value of a generic observable $A$, $\langle A\rangle_{\rho}:=\mathrm{tr}\left(A\rho\right)$, under the evolution generated by $UHU^{\dagger}$. In the Supplemental Material, Section IV, we show, assuming all windows centered around an extensive energy contain exponentially many eigenvalues, that except with probability $(N/\delta)^{2}\mathrm{e}^{-\Omega(N)}$, the time-evolution is bounded by
$|\langle A\rangle_{\rho^{UHU^{\dagger}}(t)}-\langle A\rangle_{\rho_{\infty}^{UHU^{\dagger}}}|\leq\mathrm{e}^{-\Omega(N)}+R(t),$ (18)
@@
Several natural open questions remain. Importantly, the problem of identifying precise properties of the Hamiltonian achieving a lower than worst case BE error. For these Hamiltonians, our result predicts that a vast ensemble of other Hamiltonians exists that all have indistinguishable short-time dynamics and which eventually thermalize. Furthermore, while our randomization achieves thermalization, more work may be needed about exploring the precise physical mechanisms that can be held responsible for them. On the one hand, being able to implement such transformations in a controlled way might have implications in the form of algorithms for Gibbs state preparation. On the other hand, understanding under what condition an uncertainty in either state or Hamiltonian preparation can translate into a random energy-preserving perturbation such as the one we defined might shed light on the stability of many-body localization under realistic conditions. It is the hope that this work contributes to understanding that under a “trembling hand”, most systems follow the laws of quantum statistical mechanics.